A ball is shot from the top of a building with an initial velocity of 18 m/s at an angle θ = 47 ∘ above the horizontal.What are the x and y components of the initial velocity?If a nearby building is the same height and 55 m away, how far below the top of the building will the ball strike the nearby building?Express your answer to two significant figures and include the appropriate units.
Answer #1

a)

vx = 18 cos 47 = 12.276 m/s

vy = 18 sin 47 = 13.164 m/s

======

b)

time taken by ball to reach the building

t = 55 / vx = 4.48 s

vertical displacement

h = - vy t + 0.5* gt^2

h = - 13.164* 4.48 + 4.9* 4.48^2

h = 39.37 m

=======

Comment before rate in case any doubt, will reply for sure.. goodluck